COMPUTATION

1. An inelastic collision occurs in one dimension, in which a 15 kg block traveling at 7.6 m/s collides with a 13.2 kg block traveling at 2.5 m/s in the same direction, and they stick together. What are the velocities of the blocks immediately after the collision?



1
Expert's answer
2022-06-06T14:58:14-0400

The law of conservation of momentum says

"m_1v_1+m_2v_2=(m_1+m_2)v"

"v=\frac{m_1v_1+m_2v_2}{m_1+m_2}"

"v=\frac{15*7.6+13.2*2.5}{15+13.2}=5.21\:\rm m/s"